In searching the bottom of a pool at night, a secu
Answer #1

Given,

h = 1.3 m and d = 2.5 m ; D = 2.1 m

The angle of incidence of light will be given by:

tan\theta = d/h = 2.5/1.3 = 1.923 => \theta = tan -1 ( 1.923)

\theta = 62.53 degrees.

As the light is travelling from the air into water, we can calculate the angle of refraction using Snell's law.

n1 sin\theta1 = n2 sin\theta2

1 x sin62.53 = 1.33 x sin \theta2

\theta2 = 41.84 degrees

Now the horizontal distance from the edge of the pool will be :

L = d + D tan\theta2 = 2.5 + 2.1 x tan41.84 = 4.38 m

Hence, L = 4.38 m.